OpenStudy (anonymous):

Absolute Value Equation 6 – 3|2x + 6| = 0

OpenStudy (anonymous):

6 – 3|2x + 6| = 0 - 3|2x + 6| = -6 divide both sides by -3 |2x+6| = (-6/-3) |2x+6| = 2 now you have the absolute value sign. this means that their are two possible answers. 2x + 6 = 2 AND 2x + 6 = -2 solve both equations and you get two answers, which is what you are supposed to get.